OpenStudy (marialupossalazar):

A horse-drawn carriage travels 42 2/3 kilometers in 4 1/3 hours. At this rate, how many kilometers does it travel per hour?

jabez177 (jabez177):

\[42\frac{ 2 }{ 3 }\div4\frac{ 1 }{ 3 }\]Simple.

jabez177 (jabez177):

Do you know how to divide fractions?

OpenStudy (marialupossalazar):

no

jabez177 (jabez177):

Alright then! Let's go from step 1! :)

jabez177 (jabez177):

First step is we have to convert the mixed fraction to an improper fraction.

OpenStudy (marialupossalazar):

ok

jabez177 (jabez177):



jabez177 (jabez177):

So in that example, we do 3 x 4 then + 1 which is 13 and we keep the same denominator which is 3 so we get 13/3

jabez177 (jabez177):

Do you get that?

OpenStudy (marialupossalazar):

yes thank you

jabez177 (jabez177):

So now we have to do the same to the 2 mixed fractions. :) 

jabez177 (jabez177):

3 x 42 + 2 = ?